OpenStudy (anonymous):

Write as a product. c^2+d-d^2+c

OpenStudy (anonymous):

@Callisto

hartnn (hartnn):

do i know the difference of squares formula ? \(a^2-b^2=(...+...)(...-...)\)

hartnn (hartnn):

*do you know...

OpenStudy (anonymous):

yes

hartnn (hartnn):

so what will be c^2-d^2= >?

hartnn (hartnn):

*c^2-d^2= ..... ?

hartnn (hartnn):

anybody here ?